OpenStudy (readergirl12):

what is the value of x? 4(x+2)=14-2(3-4x) x=?? HALPPPPP

OpenStudy (misty1212):

lol HAALLLPPPPP!!

OpenStudy (misty1212):

hi (again) ready?

OpenStudy (readergirl12):

Mhm.

OpenStudy (anonymous):

Just make sure you distribute and this should be very easy

OpenStudy (misty1212):

we have to remove the parentheses using the distributive property first

OpenStudy (anonymous):

yeah that ^ You got this ill leave cx

OpenStudy (smilie2015):

4(x+2) = 14-2(3-4x) 4x+8 =14 – 6 -8x 4x +8 =8 – 8x 4x +8x = 8-8 12x = 0 x=0